if you deposit 4500 in an account that earns 3.5% annual interest, whats the balance after one year if the interest is compounded annually? also monthly and quarterly?
Answer by rfer(16322)

A=4500(1+0.035)
A=4657.50
---------------
A=4500(1+0.035/12)^12
A=4500(1.0029)^12
A=4500*1.03536
A=4659.12
---------
You do the last one.